OpenStudy (anonymous):

What is an equation of a line (in point-slope form) that passes through the points (5,1) and (0,2)?

OpenStudy (abb0t):

Point-slope form: y − b = m(x − a) for a point (a, b) and slope m Slope-intercept form: y = mx + b, for slope m and y-intercept b

hero (hero):

1. Write y = mx + b 2. Insert the given points into y = mx + b 3. Isolate b in both equations 4. Solve for m, then solve for b. y = mx + b 2 = 0m + b 1 = 5m + b 2 = b 1 - 5m = b b = b 1 - 5m = 2 1 - 2 = 5m -1 = 5m -1/5 = m 2 = b Therefore, the slope-intercept form of the line is: y = -(1/5)x + 2

OpenStudy (anonymous):

thankss :))